OpenStudy (anonymous):

f f(x) = x + 5, what does f(8) represent?

OpenStudy (anonymous):

The value if x when (x + 5) = 8 The value of 8f(x) when x = 8 The value of f(x) when (x + 5) = 8 The value of (x + 5) when x = 8

OpenStudy (anonymous):

lets start with an example, if f(x)=x f(1) essentially means put x=1 into the right side of the equation so if f(x)=x f(1)=1(x=1), hope you understand joseph!

OpenStudy (anonymous):

so basically its d?

OpenStudy (anonymous):

yeap that's right!

OpenStudy (anonymous):

thanks!